# Accurate and thermodynamically consistent hydrogen equation of state for planetary modeling with flow matching

### v1

@@ Abstract
Introduction. The equation of state (EOS) of dense hydrogen (H) is one of the core ingredients to model the internal structure of giant gaseous planets, such as Jupiter and Saturn [1]. With the wealth of data from recent spacecraft missions, Galileo [2] and Juno [3], which measured Jupiter’s atmospheric composition and gravitational field, lack of accurate knowledge of the thermodynamic properties of hydrogen is now considered as one of the major roadblocks in planetary science [4, 5, 6, 7]. As parameterized planetary models already require several hypotheses [5, 8, 9, 10], an uncontested EOS is needed to advance the field.
@@
Since the high temperature and pressure inside these planets are still largely beyond the reach of current experimental techniques, one usually relies instead on numerical simulations, typically ab initio molecular dynamics (MD) based on electron structure methods such as density functional theory (DFT) or quantum Monte Carlo (QMC). Up to now, there have been several widely used ab initio EOSs for planetary science applications, including the one by Chabrier, Mazevet and Soubiran (CMS19) [11], by Militzer and Hubbard (MH13) [12], and the Rostock EOS (REOS2 [13], REOS3 [14]) of various versions. All of them are derived from DFT-MD simulations using the exchange-correlation functional of Perdew, Burke and Ernzerhof (PBE) [15]. Despite this common origin, however, they can make significant differences in the resulting Jupiter model, such as the location of adiabat and the predicted size of core mass [4, 7].
@@
The key factor behind these discrepancies of planetary models is the calculation of entropy, which has been recognized and discussed to some extent in several studies [12, 4, 6]. Unlike observables such as energy and pressure , the Helmholtz free energy and entropy are not directly accessible from ab initio simulations and often rely on extra procedures called thermodynamic integration (TI). One common implementation of this technique involves interpolating the available data to create a continuous path on the temperature-density phase diagram [13, 4]. From basic statistical mechanics we have
@@
These two equations correspond to integrating the energy (pressure) of the system along isochores (isotherms), respectively. However, this approach has been criticized because, due to limited number of simulation points in practice, the interpolation procedure may introduce uncontrollable systematic errors [12].
@@
An alternative TI scheme is called the coupling constant integration (CCI) [16, 17] or Hamiltonian thermodynamic integration [18]. By connecting the target state to an artificial Hamiltonian with tractable thermodynamic properties, this method further allows for the calculation of absolute entropy. However, it is also much more expensive due to the need to perform additional MD simulations, and therefore only suitable for investigating a small region of the phase diagram [19, 12]. Besides, one has to carefully choose relevant parameters in the artificial system and MD simulations to ensure well-converged results [20], and the errors associated with discrete simulation points still remain.
@@
The uncertainties involved in entropy calculation are further exacerbated by the fact that the EOS used for planetary modeling is usually composed of several parts, each based on different theoretical approach and has different phase region of validity. For example, although DFT-MD can provide a good description of electron correlations at intermediate temperatures and densities, especially near pressure-induced dissociation, it will quickly become inefficient at lower densities . For such a region deep within the molecular phase, empirical chemical models such as the well known Saumon–Chabrier–van Horn (SCvH) EOS [21] are generally believed to be accurate and reliable. In practice, the data from these different methods needs to be connected in some way. This, however, could possibly introduce critical errors, as we will demonstrate below.
In this Letter, we focus on the crucial issue of entropy calculation at the DFT-PBE level of theory, aiming to resolve long-standing disagreements among various Jupiter model predictions, which are highly sensitive to small changes in entropy [12, 4, 22, 6]. We rely solely on ab initio theory and well-established experimental data without invoking indirect observations. The purpose of this paper is not to reconcile planetary models with spacecraft measurements [1, 5], but rather to provide a non-parametric hydrogen EOS, ranging from roughly 1 bar to 700GPa, that forms a solid foundation for future model revisions.
We calculate the free energy and entropy of dense hydrogen using a recently developed method called flow matching [23, 24, 25]. Unlike the demonstrative example shown in Ref. [26], here the calculations are based on DFT-MD simulations and thus have ab initio accuracy. Basically, we employ a machine-learned invertible transformation , which acts on configuration samples of the proton coordinates, to directly connect two phase points of interest. Given the transformation can bring the two states close enough to ensure a sufficient overlap, their free energy difference (scaled by temperature) can then be efficiently estimated as follows:
where are the corresponding Boltzmann distributions, , being the Boltzmann constant, and
are known as the forward and reverse work, respectively [27]. In practice, the two estimators in Eq. (2) provide both upper and lower bound of the free energy difference, which can be pretty tight [27]. In this way, one can accurately estimate the entropy difference based solely on samples and energy functions of the two end states, thus avoid errors arising from any kinds of interpolation between them as done in TI. Moreover, it is also fairly straightforward to compute the absolute entropy by just choosing one of the state to have tractable thermodynamic properties.
The computational framework of Eqs. (2) and (3) is usually known as targeted free energy perturbation (TFEP) [28, 29]. Thanks to the rapid development of deep generative models in recent years, the desired invertible map can now be efficiently represented by a class of generative models called normalizing flow [30], making TFEP emerge as a promising tool for free energy calculation [31, 32, 33, 34]. In contrast to the “variational” approach, which has previously been applied to a broad range of many-body problems [35, 36, 32, 33, 37, 38, 39], the flow matching method we employ in this work represents a different training scheme of the flow model [23, 24, 25], which offers a variety of advantages in both accuracy and efficiency; see Supplemental Material [27] for details.
@@
Entropy from ab initio data. We employ the flow matching method to study the ab initio portion of the phase diagram, with K and . Specifically, we repeatedly train a transformation and use Eq. (2) to compute the entropy difference for each neighboring pair of states on a grid; see Supplemental Material [27] for an example. We also calculate the absolute entropy of a reference state within the same framework and then shift the entropy of other states rigidly. In practice, we choose the reference point at and [27], where the computed value of the absolute entropy is in excellent agreement with those reported by Morales et al. [19] based on coupled electron-ion Monte Carlo (CEIMC) simulations.
@@
Figure 1 shows our results (solid line) for several isotherms of the specific entropy and pressure , together with two widely used hydrogen EOSs in past literature: CMS19 [11] (dotted line) and REOS3 [14] (dashed line). For convenience, we also include an error bar of absolute entropy at the reference point, and the CEIMC entropy of Morales et al. [19] (dash-dotted line) at relatively high densities. Note the REOS3 entropy is calculated by us using TI (see Eq. (1)), and, for sake of visualization, we have introduced a global entropy offset to make it match our data at the bottom-left phase point, i.e., and . We do not compare with the MH13 EOS [12] since it is derived directly for hydrogen-helium mixtures.
@@
One striking feature of Fig. 1(a) is the non-monotonic behavior of entropy at low temperatures in both our results and REOS3, which is qualitatively different from CMS19. Moreover, CMS19 does not feature the typical slope change of the relation, as shown in Fig. 1(b), which is emanating from the critical point of the liquid-liquid phase transition between molecular and atomic phase [40, 41]. This indicates that although derived from DFT-PBE data in the phase region considered here [11, 42, 43, 19], the CMS19 EOS does not closely capture the pressure-induced molecular dissociation, which could have a notable impact on Jupiter models [27].
@@
The observed behavior of entropy and pressure in our results, as described above, can be understood by virtue of the thermodynamic relation [21]: . In fact, by examining the pressure isochores as illustrated in the inset of Fig. 1(b), the region where the pressure decreases with temperature corresponds precisely to the region of increasing entropy in Fig. 1(a). Physically, they are just different manifestations of the molecular dissociation effects: the attractive force from broken molecular bonds can dominates the repulsion from kinetic motion, while the larger number of configurations involving only individual H atoms (compared to those with bound H2-like pairs) can compensate for the smaller occupied volume caused by increasing density.
@@
Benchmarking thermodynamic integration. We can use the flow matching entropy to benchmark the TI approach, which simply amounts to integrating Eq. (1) using our pressure and energy data. We find that the TI entropies are virtually indistinguishable from the solid lines in Fig. 1(a). Besides, they do not depend on the choice of integration path and hence show very good thermodynamic consistency. Contrary to the previous belief [12], our observations indicate that TI performs quite well for a discrete grid with spacings in the order of K and . When combined with the absolute entropy calculation on only a single referent point, as done above, we can then achieve the same level of accuracy as the “full” CCI approach in Ref. [12], but with far less computational costs.
@@
Note that compared to REOS3, we have added some additional simulation points at relatively low densities to improve the characterization of molecular dissociation. Nevertheless, TI is not fundamentally flawed even with the relatively sparse grid of REOS3 in the ab initio region, and the resulting entropy remains generally consistent with our data, as shown by the dashed lines in Fig. 1(a). However, TI becomes truly problematic when we further extend the integration path to touch other regions based on inconsistent theories. We can quantitatively illustrate such inconsistency by performing the line integral along each closed local square loop of the tabular data, which should in principle vanish everywhere. However, for REOS3 we found large deviation from this ideal behavior between the ab initio region and the chemical model at lower densities [44, 45], as shown in Figure 2. This implies that the resulting entropy will be ambiguously defined and strongly depend on the choice of integration path and initial reference point; see Supplemental Material [27] for more details. One can then identify an entropy error of approximately [27], which yields qualitatively large differences on planetary models derived in the past using this method [13, 4].
@@
A thermodynamically consistent construction. We have shown that performing a single TI over an EOS containing various inconsistent theories can lead to significant and uncontrolled errors of the entropy. A better approach is to perform TI on each component (i.e., region where the same theory applies) of the EOS separately, and then interpolate the resulting free energy function [27]. In this way, one can derive the pressure and energy near the interpolation boundaries by taking partial derivatives, which are naturally consistent by construction. Another advantage is that one can easily incorporate accurate values of the absolute entropy into each separate region. This ensures that any residual interpolation errors that one can make at the boundaries remain localized and do not propagate elsewhere.
@@
Figure 3(a-d) show our final EOS (solid lines), which consist of the ab initio DFT-MD part at , the SCvH EOS at , and an interpolation region between them. For comparison, we also include the entropy and pressure isotherms of Miguel et al. [4] (dashed lines), which are obtained by a single TI over the entire phase region of REOS3. Note the new interpolation procedure depends very sensitively on the entropy difference between SCvH and the ab initio data [27]. In doing this, we then surprisingly found the reported (absolute) entropies of SCvH at room conditions (e.g., at standard pressure 1 bar and temperature 298.15K) are about (or ) higher than both the experimental data [46, 47] and theoretical calculations [48]. Only after such errors are fixed (by applying a global entropy shift on the SCvH side) can we achieve a high-quality interpolation that respects fundamental thermodynamic constraints; see Supplemental Material [27] for details. This observation, which results from totally independent experimental or numerical benchmarks at both low ( 1 bar) and high ( 700GPa) pressures, is a very strong evidence on the accuracy of our construction. Fig. 3(e) also shows the local loop integral values of our final EOS; compared to Fig. 2, one clearly sees the new interpolation approach can indeed yield much better thermodynamic consistency over the entire phase diagram, as discussed above.
@@
Jupiter’s adiabat. We demonstrate the utility of the newly constructed hydrogen EOS by performing a preliminary calculation of Jupiter’s adiabat. We assume a homogeneous model with no compact core, and set a helium mass fraction of to ensure a fair comparison with previous works [4, 22, 12]. We adopt a simple linear mixing approximation for the entropy of mixtures, using the REOS3 EOS of helium derived by Miguel et al. in Ref. [22]. The starting point is bar and K [1], which corresponds to an entropy (for hydrogen) of .
@@
Figure 4 shows our result for the adiabat (red dashed line), together with those derived from the REOS3 (black line) [22] and MH13 EOS (blue line) [12], respectively. For a comparison also with CMS19, see the Supplemental Material [27]. Note the hydrogen EOS used in the REOS3 adiabat was produced by Miguel et al. [4, 22] using a single TI over the entire phase region, as already shown by the dashed lines in Fig. 3. Our result clearly agrees well with the REOS3 adiabat up to about 1GPa, which is as expected since both the underlying hydrogen EOSs are derived at low pressures from the same SCvH model. However, at higher pressures, our adiabat begins to diverge from the REOS3 line and becomes cooler. This is a direct consequence of the systematically lower REOS3 entropy in the ab initio region in Fig. 3(a), which is in turn caused by the thermodynamic inconsistency shown in Fig. 2 [27]. Notice such a cooler adiabat passes very close to the miscibility boundary for hydrogen-helium mixtures, shown as the yellow line in Fig. 4, albeit derived from DFT-MD simulations using an exchange-correlation functional different from PBE [49]. This suggests an increased likelihood of phase separation between hydrogen and helium in Jupiter’s deep interior, causing the formation of so-called helium rain [1, 50, 51].
@@
On the other hand, our adiabat at high pressures is apparently closer to the MH13 prediction, as shown by the blue line in Fig. 4. Note that the MH13 EOS is derived directly from ab initio MD simulations of a hydrogen-helium mixture and does not rely on the linear mixing approximation. Overall, we show that, at the DFT-PBE level of theory, the long-standing discrepancy between the Jupiter models of REOS3 and MH13 primarily stems from a problematic TI calculation of the REOS3 entropy that is not everywhere thermodynamically consistent. However, according to our earlier benchmark using flow matching, this definitely does not imply that TI is inherently inferior to the CCI approach for a finite number of simulation points, as speculated in Ref. [12]. Another observation from Fig. 4 is that the MH13 adiabat [12] was derived only for the ab initio region at high pressures. In fact, Ref. [12] did not really attempt a thermodynamic consistent interpolation with other regions to construct a complete EOS, which, however, is definitely needed for a thorough study of planetary models.
@@
Summary and outlook. In this Letter, we closely examine the entropy calculation of dense hydrogen by benchmarking TI with an independent flow matching method. Based on this, we then construct a new hydrogen EOS that is both accurate and thermodynamically consistent over several orders of magnitude of pressure, which is crucial for reliable and conclusive planetary model predictions. Notably, the interpolation quality at the boundaries serves as a highly sensitive and stringent test for the accuracy of the theories on both sides [27]. Indeed, the new approach has allowed us to identify and correct a significant discrepancy in SCvH relative to experimental data at low pressures, which is fairly surprising and remarkable given the chemical model’s long history of nearly 30 years.
@@
We conclusively address the long-standing issue of entropy disagreements among various popular EOSs used in planetary modeling [11, 12, 14], all of which are derived from the same DFT-PBE level of electronic structure theory. To further improve the hydrogen EOS and potentially reconcile models with observational data [5], one can upgrade the description of electronic correlation from DFT-PBE to a higher-level theory [52, 53, 54, 39]. We also note that as a generic tool, flow matching can also be used to study other rich phases of hydrogen [55, 56] or other interesting condensed matter systems [51, 57], where accurate calculation of free energy plays a significant role.

### v2

@@ Abstract
Introduction. The equation of state (EOS) of dense hydrogen (H) is one of the core ingredients to model the internal structure of giant gaseous planets, such as Jupiter and Saturn Helled et al. (2020). With the wealth of data from recent spacecraft missions, Galileo Johnson et al. (1992) and Juno Bolton et al. (2017), which measured Jupiter’s atmospheric composition and gravitational field, the lack of accurate knowledge of the thermodynamic properties of hydrogen is now considered as one of the major roadblocks in planetary science Miguel, Y. et al. (2016); Debras and Chabrier (2019); Mazevet, S. et al. (2022); Howard, S. et al. (2023). As parameterized planetary models already require several hypotheses Debras and Chabrier (2019); Nettelmann et al. (2021); Miguel, Y. et al. (2022); Militzer et al. (2022), an uncontested EOS is needed to advance the field.
@@
Since high-pressure experiments are difficult to perform and still cannot yield precise results, one usually relies instead on numerical simulations, typically ab initio molecular dynamics (MD) based on electron structure methods such as density functional theory (DFT) or quantum Monte Carlo (QMC). Up to now, there have been several widely used ab initio EOSs for planetary science applications, including the one by Chabrier, Mazevet and Soubiran (CMS19) Chabrier et al. (2019), by Militzer and Hubbard (MH13) Militzer and Hubbard (2013), and the Rostock EOS (REOS2 Nettelmann et al. (2012) and REOS3 Becker et al. (2014)) of various versions. All of them are derived from DFT-MD simulations using the exchange-correlation functional of Perdew, Burke and Ernzerhof (PBE) Perdew et al. (1996). Despite this common origin, however, they can make significant differences in the resulting Jupiter model, such as the location of adiabat and the predicted size of core mass Miguel, Y. et al. (2016); Howard, S. et al. (2023).
@@
The key factor behind these discrepancies of planetary models is the calculation of entropy, which has been recognized and discussed to some extent in several studies Militzer and Hubbard (2013); Miguel, Y. et al. (2016); Mazevet, S. et al. (2022). Unlike observables such as energy and pressure , the Helmholtz free energy and entropy are not directly accessible from ab initio simulations and often rely on extra procedures called thermodynamic integration (TI). One common implementation of this technique involves interpolating the available data to create a continuous path on the temperature-density phase diagram Nettelmann et al. (2012); Miguel, Y. et al. (2016). From basic statistical mechanics we have
@@
These two equations correspond to integrating the energy (pressure) of the system along isochores (isotherms), respectively. However, this approach has been criticized Militzer and Hubbard (2013) because, due to limited number of MD simulation points and inevitable statistical uncertainties in practice, the interpolation procedure may introduce systematic errors that are difficult to detect.
@@
An alternative TI scheme is called the coupling constant integration (CCI) Kirkwood (1935); de Wijs et al. (1998) or Hamiltonian thermodynamic integration Frenkel and Smit (2023). By connecting the target state to an artificial Hamiltonian with known thermodynamic properties, this method further allows for the calculation of absolute entropy. However, it is also much more expensive due to the need to perform additional MD simulations and therefore only suitable for investigating a small region of the phase diagram Morales et al. (2010a); Militzer and Hubbard (2013). Besides, one has to carefully choose relevant parameters in the artificial system and MD simulations to ensure well-converged results Bergermann et al. (2024), and the errors associated with discreteness and statistical fluctuation of the simulation points still remain.
@@
The uncertainties involved in entropy calculation are further exacerbated by the fact that the EOS used for planetary modeling is usually composed of several parts, each based on different theoretical approach and has different phase region of validity. For example, although DFT-MD can provide a good description of electron correlations at intermediate temperatures and densities, especially near pressure-induced dissociation, it will quickly become inefficient at lower densities . For such a region deep within the molecular phase, empirical chemical models, such as the well-known Saumon–Chabrier–van Horn (SCvH) EOS Saumon et al. (1995) are generally believed to be accurate and reliable. In practice, the data from these different methods need to be connected in some way. This, however, could possibly introduce critical errors, as we will demonstrate below.
In this Letter, we focus on the crucial issues of entropy calculation outlined above at the DFT-PBE level of theory, aiming to resolve long-standing disagreements among various Jupiter model predictions, which are highly sensitive to small changes in entropy Militzer and Hubbard (2013); Miguel, Y. et al. (2016, 2018); Mazevet, S. et al. (2022). The work is structured as follows: First, we calculate the free energy and entropy in the DFT-MD region using a recently developed method called flow matching Albergo and Vanden-Eijnden (2023); Lipman et al. (2023); Liu et al. (2023). As discussed below, such an independent approach is highly desired to benchmark the accuracy of TI and identify crucial deficiencies in previous works. Second, (and more importantly,) we devise a simple and efficient protocol to construct an EOS that can achieve excellent thermodynamic consistency across various theories and a broad range of conditions. Our procedure is primarily designed to provide a reliable and uncontested input for planetary modeling, thereby forming a solid foundation for future model refinements against spacecraft measurements and further applicability to other materials Helled et al. (2020); Debras and Chabrier (2019).
Flow matching. Consider any two given phase points and of the hydrogen system. When calculating their free energy difference using TI, one needs to create a continuous path between them by interpolation. On the other hand, flow matching is built on the framework of targeted free energy perturbation (TFEP) Jarzynski (2002); Hahn and Then (2009), which solely relies on information about the two end states 0 and 1; see the Eqs. (2) and (3) below. As a result, it can avoid any errors arising from a manual interpolation procedure and thus serves as a valuable benchmark for TI.
More specifically, within the framework of TFEP, the free energy difference (scaled by temperature) between the two states can be written as the following estimators:
Here, denotes a configuration sample of the proton coordinates, and are the corresponding Boltzmann distributions, with , being the Boltzmann constant. The quantities
are known as the forward and reverse work, respectively SM . Note that the above equations hold in principle for any invertible transformation that acts on the configuration space. In particular, when is the identity map, Eqs. (2) and (3) reduce to the usual formulas for importance sampling. Moreover, we note that by simply selecting one of the states to have known thermodynamic properties, it is also fairly straightforward to compute the absolute free energy and entropy of the system.
In practice, due to the limited number of samples, Eq. (2) can be efficiently estimated with low variance only for some carefully designed transformation , which should bring the two states of interest close enough to ensure sufficient overlap of their probability densities. In this work, we achieve this goal by representing as a class of generative neural network called normalizing flow Papamakarios et al. (2021) and training the network using the flow matching technique Albergo and Vanden-Eijnden (2023); Lipman et al. (2023); Liu et al. (2023). Note that unlike the demonstrative example shown in Ref. Zhao and Wang (2023), here the calculations are based on DFT-MD simulations and thus have ab initio accuracy. A more detailed introduction to the relevant background is provided in the Supplemental Material SM .
@@
Entropy from ab initio data. We employ the flow matching method to study the ab initio portion of the phase diagram with K and . Specifically, we repeatedly train a transformation and use Eq. (2) to compute the entropy difference for each neighboring pair of states on a grid; see Supplemental Material SM for an example. We also calculate the absolute entropy of a reference state within the same framework and then shift the entropy of other states rigidly. In practice, we choose the reference point at and SM , where the computed value of the absolute entropy is in excellent agreement with those reported by Morales et al. Morales et al. (2010a) based on coupled electron-ion Monte Carlo (CEIMC) simulations.
@@
Figure 1 shows our results (solid line) for several isotherms of the specific entropy and pressure , together with two widely used hydrogen EOSs in past literature: CMS19 Chabrier et al. (2019) (dotted line) and REOS3 Becker et al. (2014) (dashed line). For convenience, we also include an error bar of absolute entropy at the reference point and the CEIMC entropy of Morales et al. Morales et al. (2010a) (dash-dotted line) at relatively high densities. Note that the REOS3 entropy is calculated by us using TI (see Eq. (1)), and, for the sake of visualization, we have introduced a global entropy offset to make it match our data at the bottom-left phase point, i.e., and . We do not compare with the MH13 EOS Militzer and Hubbard (2013) since it was derived directly for a hydrogen-helium mixture, rather than the pure hydrogen system studied here.
@@
One striking feature of Fig. 1(a) is the nonmonotonic behavior of entropy at low temperatures in both our results and REOS3, which is qualitatively different from CMS19. Moreover, CMS19 does not feature the typical slope change of the relation, as shown in Fig. 1(b), which is emanating from the critical point of the liquid-liquid phase transition between molecular and atomic phase Lorenzen et al. (2010); Geng et al. (2019). This indicates that although derived from DFT-PBE data in the phase region considered here Chabrier et al. (2019); Caillabet et al. (2011); Holst et al. (2008); Morales et al. (2010a), the CMS19 EOS does not closely capture the pressure-induced molecular dissociation, which could have a notable impact on Jupiter models SM .
@@
According to the thermodynamic relation Saumon et al. (1995) , the region of “abnormally” increasing entropy in Fig. 1(a) corresponds precisely to that of decreasing pressure with temperature, as illustrated by the isochores in the inset of Fig. 1(b). Physically, they are just different manifestations of the molecular dissociation effects: The attractive force from broken molecular bonds can dominate the repulsion from kinetic motion, while the larger number of configurations involving only individual H atoms (compared to those with bound H2-like pairs) can compensate for the smaller occupied volume caused by increasing density.
@@
Benchmarking thermodynamic integration. We use the flow matching entropy in the ab initio region to benchmark the TI approach, which simply amounts to integrating Eq. (1) using our pressure and energy data. We find that the TI entropies are virtually indistinguishable from the solid lines in Fig. 1(a). Besides, they do not depend on the choice of integration path and hence show very good thermodynamic consistency. Our observations indicate that TI performs quite well for a discrete grid with spacings on the order of K and . When combined with the absolute entropy calculation on only a single reference point, as done above, we can then achieve the same level of accuracy as the “full” CCI approach in Ref. Militzer and Hubbard (2013) but with far less computational costs.
@@
Note that, compared to REOS3, we have selected a slightly denser grid of simulation points between and to improve the characterization of molecular dissociation. Nevertheless, the “discreteness” error of TI in the ab initio region of REOS3 is still under control, and the resulting entropy remains generally consistent with our data, as shown by the dashed lines in Fig. 1(a). However, TI becomes truly problematic when we further extend the integration path to touch other regions based on inconsistent theories. We can quantitatively illustrate such inconsistency by performing the line integral along each closed local square loop of the tabular data, which should, in principle, vanish everywhere. However, for REOS3, we found large deviation from this ideal behavior between the ab initio region and the chemical model at lower densities Juranek et al. (2002); Holst et al. (2007), as shown in Fig. 2. This implies that the resulting entropy will be ambiguously defined and strongly depend on the choice of integration path and initial reference point; see Supplemental Material SM for more details. One can then identify an entropy error of approximately SM , which yields qualitatively large differences on planetary models derived in the past using this method Nettelmann et al. (2012); Miguel, Y. et al. (2016).
@@
Thermodynamically consistent construction. An inherent difficulty of the “single TI over the entire phase diagram” approach stems from the fact that energy and pressure are different partial derivatives of the same free energy function and hence not independent. Consequently, it turns out to be highly challenging to simultaneously interpolate both energy and pressure data produced by different theories in a consistent way, as demonstrated in Fig. 2 for the case of REOS3. A better approach is to perform TI on each component (i.e., region where the same theory applies) of the EOS separately and then interpolate the resulting free energy function SM . In this way, one can derive the pressure and energy near the interpolation boundaries by taking partial derivatives, which are naturally consistent by construction. Another advantage is that one can easily incorporate accurate values of the absolute entropy into each separate region. This ensures that any residual interpolation errors that one can make at the boundaries remain localized and do not propagate elsewhere.
@@
Figures 3(a-d) show our final EOS (solid lines), which consists of the ab initio DFT-MD part at (only above K), the SCvH EOS at (covering a broader temperature range down to K), and an interpolation region between them. For comparison, we also include the entropy and pressure isotherms of Miguel et al. Miguel, Y. et al. (2016) (dashed lines), which are obtained by a single TI over the entire phase region of REOS3. Note that the new interpolation procedure depends very sensitively on the entropy difference between SCvH and the ab initio data SM . In particular, a high-quality interpolation can only be achieved after globally subtracting a constant of from the originally reported SCvH entropy. Remarkably, this observation aligns precisely with various independent numerical Chabrier et al. (2019) and experimental references Cox et al. (1989); Lemmon et al. (2023). There, the entropy constant was recognized as , which accounts for the contribution of proton spin degrees of freedom that are implicitly neglected on the DFT-MD side Saumon and Chabrier ; see Supplemental Material SM for details. This strongly indicates that our new protocol is indeed the very approach to construct accurate and uncontested hydrogen EOS for planetary modeling. Figure 3(e) also shows the local loop integral values of our final EOS; compared to Fig. 2, one clearly sees that the new protocol can indeed yield much better thermodynamic consistency over the entire phase diagram, as discussed above.
@@
Jupiter’s adiabat. We demonstrate the utility of the newly constructed hydrogen EOS by performing a preliminary calculation of Jupiter’s adiabat. We assume a homogeneous model with no compact core and set a helium mass fraction of to ensure a fair comparison with previous works Miguel, Y. et al. (2016, 2018); Militzer and Hubbard (2013). We adopt a simple linear mixing approximation for the entropy of mixtures, using the REOS3 EOS of helium derived by Miguel et al. in Ref. Miguel, Y. et al. (2018).
@@
Figure 4 shows our result for the adiabat (red dashed line) together with those derived from the REOS3 (black line) Miguel, Y. et al. (2018) and MH13 EOS (blue line) Militzer and Hubbard (2013), respectively. For a comparison also with CMS19, see the Supplemental Material SM . Note that the hydrogen EOS used in the REOS3 adiabat was produced by Miguel et al. Miguel, Y. et al. (2016, 2018) using a single TI over the entire phase region, as already shown by the dashed lines in Fig. 3. Our result clearly agrees well with the REOS3 adiabat up to about 1GPa, which is as expected since both the underlying hydrogen EOSs are derived at low pressures from the same SCvH model. However, at higher pressures, our adiabat begins to diverge from the REOS3 line and becomes cooler. This is a direct consequence of the systematically lower REOS3 entropy in the ab initio region in Fig. 3(a), which is in turn caused by the thermodynamic inconsistency shown in Fig. 2 SM . Notice that such a cooler adiabat passes very close to the miscibility boundary for hydrogen-helium mixtures, shown by the yellow line in Fig. 4, albeit derived from DFT-MD simulations using an exchange-correlation functional different from PBE Schöttler and Redmer (2018). This suggests an increased likelihood of phase separation between hydrogen and helium in Jupiter’s deep interior, causing the formation of so-called helium rain Helled et al. (2020); Vorberger et al. (2007); Chang et al. (2024).
@@
On the other hand, our adiabat at high pressures is apparently closer to the MH13 prediction, as shown by the blue line in Fig. 4. Note that the MH13 EOS is derived directly from ab initio MD simulations of a hydrogen-helium mixture and does not rely on the linear mixing approximation. Overall, we show that, at the DFT-PBE level of theory, the long-standing discrepancy between the Jupiter models of REOS3 and MH13 primarily stems from a problematic TI calculation of the REOS3 entropy that is not everywhere thermodynamically consistent. However, according to our earlier benchmark using flow matching, this definitely does not imply that TI is inherently inferior to the CCI approach for a finite number of simulation points, as speculated in Ref. Militzer and Hubbard (2013). Another observation from Fig. 4 is that the MH13 adiabat Militzer and Hubbard (2013) was derived only for the ab initio region at high pressures. In fact, Ref. Militzer and Hubbard (2013) did not really attempt a thermodynamic consistent interpolation with other regions to construct a complete EOS, which, however, is definitely needed for a thorough study of planetary models.
@@
Summary and outlook. In this Letter, we closely examine the entropy calculation of dense hydrogen in the ab initio region by benchmarking TI with an independent flow matching method. Building on this, we then construct a new hydrogen EOS that is both accurate and thermodynamically consistent over several orders of magnitude of pressure (roughly 1bar 700GPa). Remarkably, the interpolation quality at the boundaries serves as a highly sensitive and stringent “detector” for the accuracy of the theories on both sides SM . Such a “self-correcting” mechanism of the new protocol is exactly what we need to ensure steady convergence toward a truly reliable and conclusive hydrogen EOS for planetary modeling.
@@
We conclusively address the long-standing issue of entropy disagreements among various popular EOSs used in planetary modeling Chabrier et al. (2019); Militzer and Hubbard (2013); Becker et al. (2014), all of which are derived from the same DFT-PBE level of electronic structure theory. To further improve the hydrogen EOS and potentially reconcile models with observational data Debras and Chabrier (2019), one can upgrade the description of electronic correlation from DFT-PBE to a higher-level theory Pierleoni et al. (2016); Mazzola et al. (2018); Nakano et al. (2020); Xie et al. (2023a). We also note that as a generic tool, flow matching can also be used to study other rich phases of hydrogen Morales et al. (2010b); Niu et al. (2023) or other interesting condensed matter systems Chang et al. (2024); Wang et al. (2023), where accurate calculation of free energy plays a significant role.